# Sensing atomic superfluid rotation beyond the standard quantum limit

### v1

@@ III Rotational sensing using squeezed light
Note that the spectrum is normalized and the additional term $S_{\rm{add}}$ is present only for squeezed input. Figure 4 shows the variation of the spectral density $S^{\text{out}}_{P,P}(\omega)$ with input power $\mathcal{P}_{\rm in}$, at the optimal frequency $\omega=\omega_{\rm opt}$. The figure also shows the contribution due to optical shot noise $S_{\rm{sn}}$ and radiation pressure $S_{\rm{rp}}$ (solid blue and orange lines respectively). In the absence of squeezing, the minimum of $S^{\text{out}}_{P,P}(\omega)$ at temperature $T_{c,d}=$20 nK and frequency $\omega_{\rm opt}$ gives us the minimum noise $S_{\textrm{min}}$. This is achieved at optimal input power $\mathcal{P}_{\rm in}^{\rm out}$, as shown in Fig. 4 (solid black curve). Now, if $T=0$, the noise $S_{\rm th}$ can be further lowered to ultimately achieve the standard quantum limit $S_{\textrm{SQL}}$. Using squeezed input, with amplitude $r=2$ and $\theta=\pi$, the output spectral density $S^{\text{out}}_{P,P}$ can be lowered significantly below the minimum noise, thus providing a clear advantage over unsqueezed light for lower values of $\mathcal{P}_{\rm in}$. However, the minimal noise spectral density can only get close to SQL but not below it, regardless of the value of $r$ or $\theta$. In fact, for $\theta=0$, the noise is higher than the minimum noise. This is primarily due to the measurement backaction noise becoming significant at higher powers. As such, using squeezing alone in the input optical light with OAM, the noise cannot be lowered beyond the SQL. To break the SQL one needs to counter the backaction noise using quantum nondemolition measurement [35, 36], which can be achieved with or without using squeezed light [37].
@@ IV Rotational sensing using Backaction evasion
The measurement of a quantum system typically involves its interaction with an additional detection or measuring device. During the measurement of an observable $\hat{X}$ of the system, system-detector interactions can introduce noise in the system via the variable conjugate to $\hat{X}$, namely $\hat{Y}$, which adversely affects any future measurements of the observable. This noise is called the measurement backaction noise, and was initially evaded by making an effective joint quadrature measurement of the system [36, 55]. This technique enables the possibility of conducting the quantum non-demolition measurement [56], where measurement does not affect the observed system. The use of backaction evasion (BAE) was used to overcome the SQL in optomechanics [57], with the technique later being expanded for two-mode measurements [58, 59]. Further theoretical development has focused on BAE schemes beyond the rotating wave approximation (RWA) to include the effect of counterrotating terms [46, 47, 48].
For backaction evasion, a bichromatic laser beam with two equal frequencies, but oppositely detuned from the cavity frequency, is injected into the cavity [45] that contains the BEC in a ring trap. The optical fields oscillate with amplitudes $\varepsilon_{\pm}$ and at frequencies $\omega_{\mp}=\omega_{a}\pm\delta$, with detuning $\delta=\omega_{m}=(\omega_{c}+\omega_{d})/2$ being the average frequency of the two atomic sideband modes. The Hamiltonian is given by,
@@
where we have ignored the terms due to collisions as they only cause a relative shift in the frequencies and have no significant impact on sensitivity. Moving to the driving frame $a=e^{-i\omega_{a}t}\left[\bar{a}\left(e^{-i\delta t}+e^{i\delta t}\right)+\delta a\right]$, where $\bar{a}$ and $\delta a$ are the steady state and fluctuations of the cavity field and $\delta$ is the detuning. The linearization of the operator $a$ around a monostable steady state is shown in Appendix C. An effective Hamiltonian $H_{\rm{eff}}$ is then obtained, which is given by
@@
To obtain the solution for the time-dependent equations of motion beyond the RWA, we consider an approach [46, 47] based on the Floquet expansion of the dynamical variables. The formalism is valid even in the bad cavity regime, i.e. for $\omega_{c(d)}\ll\kappa$ as the couplings are also small $G\ll\kappa$. Applying the Floquet expansion to each dynamical variable $Z(t)$ as $Z(t)=\sum_{-\infty}^{\infty}Z^{(n)}(t)e^{in\delta t}$, the variables can be transformed to the frequency domain, by applying the Fourier transform of the above equations. Equating the terms with $e^{in\delta t}$ and considering the noise terms to be stationary ($n=0$), the following expressions for the quadratures fluctuations is obtained,
@@
$\chi_{a}^{-1}(\omega-n\delta)a^{(n)}(\omega)$ $=$ $-\delta_{n,0}\sqrt{\kappa}~{}a_{in}(\omega)-iG\bar{a}\left(X_{c}^{(n-1)}+X_{c}^{(n+1)}+X_{d}^{(n-1)}+X_{d}^{(n+1)}\right)$ (30)
@@
$\delta P^{(0)}(\omega)$ $=$ $X_{a}(\omega)G\bar{a}\sqrt{2}i\operatorname{Ga}\bar{a}X_{a}(\omega)\sqrt{\kappa}~{}Q_{\rm in}^{(0)}(\omega)\cdot\big{[}\chi_{c}(\omega-\delta)+\chi_{c}(\omega+\delta)+\chi_{d}(\omega-\delta)+\chi_{d}(\omega+\delta)\big{]},$ (31)
$\delta P^{(1)}(\omega)$ $=$ $\chi_{a}(\omega-\delta)G\bar{a}\sqrt{2\kappa}~{}\cdot\big{[}\chi_{c}(\omega)\left(\omega_{m}+\Omega\right)\epsilon_{c,\rm in}^{(0)}(\omega)+\chi_{d}(\omega)\left(\omega_{m}-\Omega\right)\epsilon_{d\rm in}^{(0)}(\omega)\big{]},$ (32)
$\delta P^{(-1)}(\omega)$ $=$ $\chi_{a}(\omega+\delta)G\bar{a}\sqrt{2\kappa}\cdot\big{[}\chi_{c}(\omega)\left(\omega_{m}+\Omega\right)\epsilon_{c\rm in}^{(0)}+\chi_{d}(\omega)\left(\omega_{m}-\Omega\right)\epsilon_{d\rm in}^{(0)}(\omega)\big{]},$ (33)
@@
Now to obtain the output spectrum, the input-output formalism is again used. Let $\{A(\omega),B(\omega),C(\omega),D(\omega)\}$ be the frequency-dependent coefficients of the noise quadrature operators $Q_{\text{in}},P_{\text{in}},\epsilon_{c,\text{in}}$ and $\epsilon_{d,\text{in}}$, respectively in Eqs. (31)-(36). The expressions for these coefficients are shown in Appendix D. The homodyne detection angle is set at $\phi=\pi/2$, as the noise spectrum is analysed using the noise quadrature $P_{\rm out}(\omega)$, which is given by
@@
$+$ $C^{(n)}(\omega)\epsilon_{c\rm in}(\omega)+D^{(n)}(\omega)\epsilon_{d\rm in}$
@@
$+D^{1}(\omega-\delta)D^{-1}(-\omega+\delta)\chi_{dd}(\omega-\delta)$
@@
Figure 6(a) shows the output spectral function $S_{P,P}^{\rm out}(\omega)$ for different values of the input power $\mathcal{P}_{\textrm{in}}=\mathcal{P}_{\pm}$, with the peak occurring at $\omega=\Omega$. The peak is directly related to the winding number $L_{p}$ of the BEC and is sharper for higher input power imbalance. The measurement sensitivity is given by $\zeta(\omega)$ defined in Eq. (13), where the optomechanical measurement time is given by $t^{-1}_{\rm{meas}}(\omega)=\big{|}{\sqrt{2\kappa}G\bar{a}}/{\left(-i\omega+\kappa/2\right)}\big{|}^{2}$. This is defined as the absolute value of the rate at which the output quadrature mode $P^{(0)}_{\rm out}$ changes with respect to the mechanical fluctuations $\delta X_{c(d)}^{(\pm 1)}$. In Fig. 6(b), the sensitivity $\zeta$ is varied with frequency $\omega$, for different input powers $\mathcal{P}_{\pm}$. Similar to the squeezed input case, the maximum sensitivity occurs at two points in the close vicinity of the spectral peak $\Omega$. The optimal frequency $\omega_{\rm opt}$ can then be used to precisely sense the angular momentum of the rotating BEC inside the cavity.
@@
The sensitivity of the backaction evasion measurements can be enhanced and the output noise lowered by adding squeezing to the bichromatic drive [49]. As in Sec. III, in this case the squeezing is introduced through the input noise correlations, with amplitude $r$ and angle $\theta$. The enhancement of sensitivity with squeezed light in the BAE method is shown in Fig. 7, for different parameter regimes. In all cases, the enhancement factor is the ratio of the optimal sensitivity, with and without squeezed light in the backaction evasion approach, varied along with some other parameter such as power or OAM of the input light.
For instance in Fig. 7(a), the enhancement factor is $\zeta_{\rm opt}(0,0)/\zeta_{\rm opt}(r,\theta)$, which shows the variation of enhancement with $r$ and $\theta$. It is observed that $\theta=\pi$ always generates the maximum sensitivity and is thus fixed for all the other plots. In Fig. 7(b), the variation of the enhancement factor with the input power $\mathcal{P}_{\rm in}$ and squeezing amplitude $r$ is shown. At low squeezing, the sensitivity decreases with higher power and the enhancement factor can increase by nearly two orders of magnitude. Importantly, at high $\mathcal{P}_{\rm in}$, where the radiation pressure noise is dominant, the effect of squeezing becomes marginal. The variation of enhancement factor with OAM number $l$ of the input light is shown in Fig. 7(c), where it is evident that enhancement is higher for larger values of $l$ and $r$. Finally, the enhancement in sensitivity is shown to only marginally change for different values of the angular momentum of the condensate ${{L_{p}\hbar}}$, as shown in Fig. 7(d).
@@ V Conclusion
By exploiting the theoretical framework of optomechanics, we have shown that the quantum sensing of such rotational properties can be significantly enhanced. The introduction of squeezed light as input, and for optimal homodyne detection angles, the sensitivity of detecting the spectral peaks of the output light and therefore the angular momentum, can be enhanced by a factor of $2-100$ ($\approx 4-20~{}{\rm dB}$) as compared to unsqueezed input. Notably, the higher enhancement occurs only at low input powers, where shot noise is dominant. Moreover, the optical quadrature noise, at optimal frequencies $\omega$, is observed to go below the shot noise and break the minimum noise level at temperature $T$, which was hitherto the best that could have been achieved using OAM light as input, without any squeezing.
@@ Acknowledgements.
Appendix A Coefficients of output quadratures
@@
Figure 10: Variation of output optical spectrum and sensitivity with detection angle $\phi$. The plots in (a)-(e) and (c)-(g), are for output spectrum and sensitivity, respectively, for unsqueezed input light, while (b)-(d) and (d)-(h) are the same, but for squeezed input with amplitude $r=2$ and angle $\theta=\pi$. The plots in the top (bottom) row correspond to cases with (without) inter-atomic interaction or collisions.
@@
$+$ $C_{P}(\omega)\epsilon_{c,\rm in}(\omega)+D_{P}(\omega)\epsilon_{d\rm in},$
@@
$+$ $C_{Q}(\omega)\epsilon_{c,\rm in}(\omega)+D_{Q}(\omega)\epsilon_{d\rm in},$
@@
Appendix B Optimal homodyne detection angle for noise spectrum and sensitivity

### v2

@@ III Rotational sensing using squeezed light
Note that the spectrum is normalized and the additional term $S_{\rm{add}}$ is present only for squeezed input. Figure 4 shows the variation of the spectral density $S^{\text{out}}_{P,P}(\omega)$ with input power $\mathcal{P}_{\rm in}$, at the optimal frequency $\omega=\omega_{\rm opt}$. The figure also shows the contribution due to optical shot noise $S_{\rm{sn}}$ and radiation pressure $S_{\rm{rp}}$ (solid blue and orange lines respectively). In the absence of squeezing, the minimum of $S^{\text{out}}_{P,P}(\omega)$ at temperature $T_{c,d}=$20 nK and frequency $\omega_{\rm opt}$ gives us the minimum noise $S_{\textrm{min}}$. This is achieved at optimal input power $\mathcal{P}_{\rm in}^{\rm out}$, as shown in Fig. 4 (solid black curve). Now, if $T_{c,d}=0$, the noise $S_{\rm th}$ can be further lowered to ultimately achieve the standard quantum limit $S_{\textrm{SQL}}$. Using squeezed input, with amplitude $r=2$ and $\theta=\pi$, the output spectral density $S^{\text{out}}_{P,P}$ can be lowered significantly below the minimum noise, thus providing a clear advantage over unsqueezed light for lower values of $\mathcal{P}_{\rm in}$. However, the minimal noise spectral density can only get close to SQL but not below it, regardless of the value of $r$ or $\theta$. In fact, for $\theta=0$, the noise is higher than the minimum noise. This is primarily due to the measurement backaction noise becoming significant at higher powers. As such, using squeezing alone in the input optical light with OAM, the noise cannot be lowered beyond the SQL. To break the SQL one needs to counter the backaction noise using quantum nondemolition measurement [37, 38], which can be achieved with or without using squeezed light [39].
@@ IV Rotational sensing using Backaction evasion
The measurement of a quantum system typically involves its interaction with an additional detection or measuring device. During the measurement of an observable $\hat{X}$ of the system, system-detector interactions can introduce noise in the system via the variable conjugate to $\hat{X}$, namely $\hat{Y}$, which adversely affects any future measurements of the observable. This noise is called the measurement backaction noise, and was initially evaded by making an effective joint quadrature measurement of the system [38, 61, 62]. This technique enables the possibility of conducting the quantum non-demolition measurement [63], where measurement does not affect the observed system. The use of backaction evasion (BAE) was used to overcome the SQL in optomechanics [64], with the technique later being expanded for two-mode measurements [65, 66]. Further theoretical development has focused on BAE schemes beyond the rotating wave approximation (RWA) to include the effect of counterrotating terms [48, 49, 50].
For backaction evasion, a bichromatic laser beam with two frequencies $\omega_{\mp}=\omega_{a}\pm\delta$ is injected into the cavity [47] that contains the BEC in a ring trap, where $\delta=\omega_{m}=(\omega_{c}+\omega_{d})/2$ is the average frequency of the two atomic sideband modes. The optical fields oscillate with amplitudes $\varepsilon_{\pm}$. The Hamiltonian is given by,
@@
where we have ignored the terms due to collisions as they only cause a relative shift in the frequencies and have no significant impact on sensitivity. Moving to the driving frame $a=e^{-i\omega_{a}t}\left[\bar{a}\left(e^{-i\delta t}+e^{i\delta t}\right)+\delta a\right]$, where $\bar{a}$ and $\delta a$ are the steady state and fluctuations of the cavity field and $\delta$ is the detuning. The linearization of the operator $a$ around a monostable steady state is shown in Appendix D. An effective Hamiltonian $H_{\rm{eff}}$ is then obtained, which is given by
@@
To obtain the solution for the time-dependent equations of motion beyond the RWA, we consider an approach [48, 49] based on the Floquet expansion of the dynamical variables. The formalism is valid even in the bad cavity regime, i.e. for $\omega_{c~{}(d)}\ll\kappa$ as the couplings are also small $G\ll\kappa$. Applying the Floquet expansion to each dynamical variable $Z(t)$ as $Z(t)=\sum_{-\infty}^{\infty}Z^{(n)}(t)e^{in\delta t}$, the variables can be transformed to the frequency domain, by applying the Fourier transform of the above equations. Equating the terms with $e^{in\delta t}$ and considering the noise terms to be stationary ($n=0$), the following expressions for the quadratures fluctuations is obtained,
@@
$\chi_{a}^{-1}(\omega-n\delta)a^{(n)}(\omega)$ $=$ $-\delta_{n,0}\sqrt{\kappa}~{}a_{\mathrm{in}}(\omega)-iG\bar{a}\left(X_{c}^{(n-1)}+X_{c}^{(n+1)}+X_{d}^{(n-1)}+X_{d}^{(n+1)}\right)$ (30)
@@
$\delta P^{(0)}(\omega)$ $=$ $\sqrt{2}\chi^{2}_{a}(\omega)G^{2}\bar{a}^{2}\kappa~{}Q_{\rm in}^{(0)}(\omega)\cdot\big{[}\chi_{c}(\omega-\delta)+\chi_{c}(\omega+\delta)+\chi_{d}(\omega-\delta)+\chi_{d}(\omega+\delta)\big{]},$ (31)
$\delta P^{(1)}(\omega)$ $=$ $\chi_{a}(\omega-\delta)G\bar{a}\sqrt{2\kappa}~{}\cdot\big{[}\chi_{c}(\omega)\left(\omega_{m}+\Omega\right)\epsilon_{c,\rm in}^{(0)}(\omega)+\chi_{d}(\omega)\left(\omega_{m}-\Omega\right)\epsilon_{d,\rm in}^{(0)}(\omega)\big{]},$ (32)
$\delta P^{(-1)}(\omega)$ $=$ $\chi_{a}(\omega+\delta)G\bar{a}\sqrt{2\kappa}\cdot\big{[}\chi_{c}(\omega)\left(\omega_{m}+\Omega\right)\epsilon_{c,\rm in}^{(0)}+\chi_{d}(\omega)\left(\omega_{m}-\Omega\right)\epsilon_{d,\rm in}^{(0)}(\omega)\big{]},$ (33)
@@
Now to obtain the output spectrum, the input-output formalism is again used. Let $\{A(\omega),B(\omega),C(\omega),D(\omega)\}$ be the frequency-dependent coefficients of the noise quadrature operators $Q_{\text{in}},P_{\text{in}},\epsilon_{c,\text{in}}$ and $\epsilon_{d,\text{in}}$, respectively in Eqs. (31)-(36). The expressions for these coefficients are shown in Appendix E. The homodyne detection angle is set at $\phi=\pi/2$, as the noise spectrum is analysed using the noise quadrature $P_{\rm out}(\omega)$, which is given by
@@
$+$ $C^{(n)}(\omega)\epsilon_{c,\rm in}(\omega)+D^{(n)}(\omega)\epsilon_{d,\rm in}$
@@
$\left.+D^{1}(\omega-\delta)D^{-1}(-\omega+\delta)\chi_{dd}(\omega-\delta)\right]$
@@
where $\chi_{i,j}(\omega)$ are the input noise correlations between $i$ and $j$ modes such that $i,j\in\{c_{\rm in}(\omega),d_{\rm in}(\omega),Q_{\rm in}(\omega),P_{\rm in}(\omega)\}$. Figure 6(a) shows the output spectral function $S_{P,P}^{\rm out}(\omega)$ for different values of the input power $\mathcal{P}_{\textrm{in}}=\mathcal{P}_{\pm}$, with the peak occurring at $\omega=\Omega$. The peak is directly related to the winding number $L_{p}$ of the BEC and is sharper for higher input power imbalance. The measurement sensitivity is given by $\zeta(\omega)$ defined in Eq. (13), where the optomechanical measurement time is given by $t^{-1}_{\rm{meas}}(\omega)=\big{|}{\sqrt{2\kappa}G\bar{a}}/{\left(-i\omega+\kappa/2\right)}\big{|}^{2}$. This is defined as the absolute value of the rate at which the output quadrature mode $P^{(0)}_{\rm out}$ changes with respect to the mechanical fluctuations $\delta X_{c(d)}^{(\pm 1)}$. In Fig. 6(b), the sensitivity $\zeta$ is varied with frequency $\omega$, for different input powers $\mathcal{P}_{\pm}$. Similar to the squeezed input case, the maximum sensitivity occurs at two points in the close vicinity of the spectral peak $\Omega$. The optimal frequency $\omega_{\rm opt}$ can then be used to precisely sense the angular momentum of the rotating BEC inside the cavity.
@@
The sensitivity of the backaction evasion measurements can be enhanced and the output noise lowered by adding squeezing to the bichromatic drive [51].
As in Sec. III, in this case the squeezing is introduced through the input noise correlations, with amplitude $r$ and angle $\theta$. The enhancement of sensitivity with squeezed light in the BAE method is shown in Fig. 7, for different parameter regimes. In all cases, the enhancement factor is the ratio of the optimal sensitivity, with and without squeezed light in the backaction evasion approach, varied along with some other parameter such as power or OAM of the input light.
For instance in Fig. 7(a), the enhancement factor is $\zeta_{\rm opt}(0,0)/\zeta_{\rm opt}(r,\theta)$, which shows the variation of enhancement with $r$ and $\theta$. It is observed that $\theta=\pi$ always generates the maximum sensitivity and is thus fixed for all the other plots. In Fig. 7(b), the variation of the enhancement factor with the input power $\mathcal{P}_{\rm in}$ and squeezing amplitude $r$ is shown. At low squeezing, the sensitivity decreases with higher power and the enhancement factor can increase by nearly two orders of magnitude. Importantly, at high $\mathcal{P}_{\rm in}$, where the radiation pressure noise is dominant, the effect of squeezing becomes marginal.
The variation of enhancement factor with OAM number $l$ of the input light is shown in Fig. 7(c), where it is evident that enhancement is higher for larger values of $l$ and $r$. Finally, the enhancement in sensitivity is shown to only marginally change for different values of the angular momentum of the condensate ${{L_{p}\hbar}}$, as shown in Fig. 7(d).
@@ V Conclusion
By exploiting the theoretical framework of optomechanics, we have shown that the quantum sensing of such rotational properties can be significantly enhanced. The introduction of squeezed light as input, and for optimal homodyne detection angles, the sensitivity of detecting the spectral peaks of the output light and therefore the angular momentum, can be enhanced by a factor of $2-100$ ($\approx 4-20~{}{\rm dB}$) as compared to unsqueezed input. Notably, the higher enhancement occurs only at low input powers, where shot noise is dominant. Moreover, the optical quadrature noise, at optimal frequencies $\omega$, is observed to go below the shot noise and break the minimum noise level at temperature $T_{c,d}$, which was hitherto the best that could have been achieved using OAM light as input, without any squeezing.
@@ Acknowledgements.
Appendix A Scattering of BEC with optical lattice formed by OAM beams
As stated in Sec. II, the cavity is injected with a single mode, symmetric superposition of OAM light given by |+=|$\ket{+}=\ket{+l\hbar}+\ket{-l\hbar}$ (with annihilation operator $a$). The light-matter interaction varies azimuthally as $\sim\cos^{2}(l\varphi)a^{\dagger}a$. This changes the spatial mode of BEC from $L_{p}$ as follows:
$\cos^{2}(l\varphi)\left|L_{p}\right\rangle$ $=$ $\left[\frac{1+\cos(2l\varphi)}{2}\right]\left|L_{p}\right\rangle,$
$=$ $\frac{1}{2}\left|L_{p}\right\rangle+\frac{1}{4}\left(e^{i2l\varphi}+e^{-i2l\varphi}\right)\left|L_{p}\right\rangle,$
$=$ $\frac{1}{2}\left|L_{p}\right\rangle+\frac{1}{4}\left(\left|L_{p}+2l\right\rangle+\left|L_{p}-2l\right\rangle\right).$
In the last line, we have used the fact that the exponential of the angular displacement operator $\varphi$ is the generator of angular momentum translations. Now the BEC order parameter in the position representation is given by
$\Phi(\varphi)$ $\sim\left\langle\varphi\left|\cos^{2}(l\varphi)\right|L_{p}\right\rangle\sim\frac{1}{2}\left\langle\varphi\mid L_{p}\right\rangle$
$+\frac{1}{4}\left(\left\langle\varphi\mid L_{p}+2l\right\rangle+\left\langle\varphi\mid L_{p}-2l\right\rangle\right)$
$\sim\frac{e^{iL_{p}\varphi}}{2}+\frac{e^{i\left(L_{p}+2l\right)\varphi}+e^{i\left(L_{p}-2l\right)\varphi}}{4}$
$\sim\frac{e^{iL_{p}\varphi}}{2}\left[1+\cos(2l\varphi)\right].$ (42)
The density is given by $n(\varphi)\sim|\Phi(\varphi)|^{2}\sim\left[1+\cos(2l\varphi)\right]^{2}$, which shows that only the $\cos(2l\varphi)$ and $\cos(4l\varphi)$ density modes are populated by scattering from the $\ket{+}$ mode.
Importantly, sine density modes, which correspond to light-matter interaction arising from antisymmetric superposition of OAM light |=|$\ket{-}=\ket{+l\hbar}-\ket{-l\hbar}$, are not present if only first-order Bragg scattering processes are considered. In fact, since higher-order scattering terms involve repetitive action of $\cos^{2}(l\varphi)$ on the scattered states, these processes also do not introduce sine mode density excitations. However, interference between the persistent currents can introduce sine modes. Nonetheless, these condensate excitations cannot transfer photons out from the $\ket{+}$ mode to any other optical mode, since Bragg scattering involves stimulated emission, which can repopulate only the $\ket{+}$ mode [67]. Photons can be scattered into modes other than $\ket{+}$ only via spontaneous emission. However, for the large light-atom detunings considered in this work, spontaneous emission is negligible [14, 26]. Thus, we only consider the $\ket{+}$ optical mode in our model.
Appendix B Coefficients of output quadratures
@@
Figure 10: Variation of output optical spectrum and sensitivity with detection angle $\phi$. The plots in (a)-(e) and (c)-(g), are for output spectrum and sensitivity, respectively, for unsqueezed input light, while (b)-(f) and (d)-(h) are the same, but for squeezed input with amplitude $r=2$ and angle $\theta=\pi$. The plots in the top (bottom) row correspond to cases with (without) inter-atomic interaction or collisions.
@@
$+$ $C_{P}(\omega)\epsilon_{c,\rm in}(\omega)+D_{P}(\omega)\epsilon_{d,\rm in},$
@@
$+$ $C_{Q}(\omega)\epsilon_{c,\rm in}(\omega)+D_{Q}(\omega)\epsilon_{d,\rm in},$
@@
Appendix C Optimal homodyne detection angle for noise spectrum and sensitivity